# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ IV Scenario for the crossover from short-range to long-range universality
Sak’s criterion $\sigma_{*}=2-\eta_{\rm SR}$ has long been regarded as a natural refinement of the original Fisher picture, ensuring a continuous function of $\eta(\sigma)$ across the LR–SR boundary. However, a closer inspection reveals that the logic behind this conjecture relies on several fragile assumptions.
First, the relation $\eta=2-\sigma$ originates from perturbative RG expansions around the Gaussian fixed point, i.e., near $\sigma=d/2$,
$\eta=2-\sigma+O(\epsilon^{3}),\,\,\,\,\,\nu=\frac{1}{\sigma}+\frac{4}{\sigma d}\frac{n+2}{n+8}\epsilon+O(\epsilon^{2})\;,$ (14)
with $\epsilon=2\sigma-d\geq 0$. Although this form remains surprisingly accurate up to three loops, there is no solid theoretical reason for it to hold as $\sigma\to 2$. Actually, in Eq. (14), the correlation-length exponent $\nu$ shows a clear dependence on parameters $\epsilon$, $n$, and $d$, indicating that the universality class cannot be fully characterized by $\eta$ alone. It remains an open question whether the perturbative result remains valid for a wide range from $\sigma=d/2$ to $2-\eta_{\rm SR}$, an assumption underlying Sak’s argument.
Second, discontinuous changes of critical exponents in parameter space are not pathological but rather expected within the RG framework. Transitions from second-order to first-order, or jumps in exponents at tricritical or multicritical points, are well-established phenomena [blumeTheoryFirstOrderMagnetic1966, capelPossibilityFirstorderPhase1966, blumeIsingModelTransition1971]. Therefore, a discontinuity of $\eta$ at $\sigma_{*}$ cannot by itself be used to dismiss the scenario $\sigma_{*}=2$.
@@
In contrast, the dominance of the $q^{\sigma}$ term over the $q^{2}$ term can be demonstrated in the Goldstone-mode physics of the LR-XY and LR-Heisenberg models in the low-T LRO phase, and can be proven for the LR-SRW (Lev́y flight) [yao2024nonclassical, 2dlrheisen]. For the LR-percolation and LR-FK-Ising model, this dominance can also be illustrated by the geometric structures of FK-random clusters at low temperatures. The crossover occurs where the analytic $q^{2}$ term overtakes the nonanalytic $q^{\sigma}$ contribution—namely at $\sigma_{*}=2$.
Taken together, these considerations demonstrate that Sak’s criterion, though elegant, rests on fragile assumptions. In contrast, a unified picture emerges from both theoretical reasoning and extensive numerical results across 2D systems – XY, Heisenberg, percolation, and Ising [yao2024nonclassical, 2dlrheisen, 2dlrperco]. In all cases, the universality class changes sharply at $\sigma=2$, where the anomalous dimension $\eta$ exhibits a weak but distinct jump. This establishes a fourth, self-consistent scenario: the exponents vary smoothly within the LR regime, yet the crossover at $\sigma_{*}=2$ is discontinuous, indicating that the boundary point already belongs to the LR universality class rather than the SR one.
A more natural argument is that, for $\sigma>d/2$, the Gaussian fixed point becomes unstable, and the interplay of the non-analytic term $q^{\sigma}$, short-range fluctuations $q^{2}$, and interaction terms would lead to rich nonclassical critical behaviors. Based on insights from the LR-SRW (i.e., Levy flights), recent high-precision Monte Carlo results, and rigorous mathematical analyses, we propose three universality diagrams of the phase transitions in the $(d,\sigma)$ plane, respectively, for the percolation, O($n$) spin, and FK-Ising models [hdlr]. For the O(n)(n) spin models, the nonclassical regime occurs in the range $d/2<\sigma\leq 2$, similarly to Ref. [fisher1972]. The dynamics of the LR-SRW changes, from diffusive for $\sigma>2$, into superdiffusive for $\sigma\leq 2$ (logarithmically super-diffusive at $\sigma=2$). Depending on $d$, this regime could be further divided into two sub-regimes separated by a new threshold $\sigma=1$. For $\sigma<1$, the dynamics of the LR-SRW are hyper-ballistic, as dominated by extremely long-range jumps that effectively short-circuit the local lattice metric. It is conjectured [hdlr] that, in sub-regime $1<\sigma\leq 2$, both $\eta$ and $\nu$ take non-trivial values (neither mean-field-like nor short-range-like), and, in sub-regime $d/2<\sigma\leq 1$, the $\nu$ value is non-trivial but $\eta=2-\sigma$ holds exactly true. For $d\geq 2$, this is consistent with the fourth scenario in Fig. 1.

### v2

@@ IV Scenario for the crossover from short-range to long-range universality
Sak’s criterion $\sigma_{*}=2-\eta_{\rm SR}$ has long been regarded as a natural refinement of the original Fisher picture, ensuring a continuous function of $\eta(\sigma)$ across the LR–SR boundary [48]. Given the formidable complexity of field theories with LR interactions, Sak’s formulation represented a substantial theoretical achievement. Nevertheless, upon closer examination, one finds that the argument ultimately rests on a set of assumptions whose validity may be limited.
First, the relation $\eta=2-\sigma$ originates from perturbative RG expansions [18] around the Gaussian fixed point, i.e., near $\sigma=d/2$,
$\eta=2-\sigma+O(\epsilon^{3}),\quad\frac{1}{\nu}=\sigma-\frac{(n+2)}{(n+8)}\epsilon+O(\epsilon^{2}),$ (14)
with $\epsilon=2\sigma-d\geq 0$. This expansion is obtained by fixing $\sigma$ and expanding in $d$. In the limit $n\to\infty$, it recovers the LR spherical model [29] result $1/\nu=d-\sigma$. However, if one fixes $d$ and expands in $\sigma$, the expression becomes
$\nu=\frac{1}{\sigma}+\frac{4}{\sigma d}\frac{(n+2)}{(n+8)}\Delta\sigma+O(\Delta\sigma^{2}),$ (15)
with $\Delta\sigma=\sigma-\frac{1}{2}d\geq 0$. This expression does not recover the LR spherical model result in the limit $n\to\infty$, revealing the subtlety of the LR perturbative expansion.
Moreover, although this form remains surprisingly accurate up to three loops for $\eta$, there is no solid theoretical reason for it to hold as $\sigma\to 2$. Actually, in Eq. (14), the correlation-length exponent $\nu$ shows a clear dependence on parameters $\epsilon$, $n$, and $d$, indicating that the universality class cannot be fully characterized by $\eta$ alone. It remains an open question whether the perturbative result remains valid for a wide range from $\sigma=d/2$ to $2-\eta_{\rm SR}$, an assumption underlying Sak’s argument.
Second, discontinuous changes of critical exponents in parameter space are not pathological but rather expected within the RG framework. Transitions from second-order to first-order, or jumps in exponents at tricritical or multicritical points, are well-established phenomena [9, 12, 8]. Therefore, a discontinuity of $\eta$ at $\sigma_{*}$ cannot by itself be used to dismiss the scenario $\sigma_{*}=2$.
@@
In contrast, the dominance of the $q^{\sigma}$ term over the $q^{2}$ term can be demonstrated in the Goldstone-mode physics of the LR-XY and LR-Heisenberg models in the low-T LRO phase, and can be proven for the LR-SRW (Lev́y flight) [60, 59]. For the LR-percolation and LR-FK-Ising model, this dominance can also be illustrated by the geometric structures of FK-random clusters at low temperatures. The crossover occurs where the analytic $q^{2}$ term overtakes the nonanalytic $q^{\sigma}$ contribution—namely at $\sigma_{*}=2$.
Taken together, these considerations demonstrate that Sak’s criterion, though elegant, rests on fragile assumptions. In contrast, a unified picture emerges from both theoretical reasoning and extensive numerical results across 2D systems – XY, Heisenberg, percolation, and Ising [60, 59, 36]. In all cases, the universality class changes sharply at $\sigma=2$, where the anomalous dimension $\eta$ exhibits a weak but distinct jump. This establishes a fourth, self-consistent scenario: the exponents vary smoothly within the LR regime, yet the crossover at $\sigma_{*}=2$ is discontinuous, indicating that the boundary point already belongs to the LR universality class rather than the SR one.
A more natural argument is that, for $\sigma>d/2$, the Gaussian fixed point becomes unstable, and the interplay of the non-analytic term $q^{\sigma}$, short-range fluctuations $q^{2}$, and interaction terms would lead to rich nonclassical critical behaviors. Based on insights from the LR-SRW (i.e., Lévy flights), recent high-precision Monte Carlo results, and rigorous mathematical analyses, we propose three universality diagrams of the phase transitions in the $(d,\sigma)$ plane, respectively, for the percolation, O($n$) spin, and FK-Ising models [53]. For the O(n)(n) spin models, the nonclassical regime occurs in the range $d/2<\sigma\leq 2$, similarly to Ref. [18]. The dynamics of the LR-SRW changes, from diffusive for $\sigma>2$, into superdiffusive for $\sigma\leq 2$ (logarithmically super-diffusive at $\sigma=2$). Depending on $d$, this regime could be further divided into two sub-regimes separated by a new threshold $\sigma=1$. For $\sigma<1$, the dynamics of the LR-SRW are hyper-ballistic, as dominated by extremely long-range jumps that effectively short-circuit the local lattice metric. It is conjectured [53] that, in sub-regime $1<\sigma\leq 2$, both $\eta$ and $\nu$ take non-trivial values (neither mean-field-like nor short-range-like), and, in sub-regime $d/2<\sigma\leq 1$, the $\nu$ value is non-trivial but $\eta=2-\sigma$ holds exactly true. For $d\geq 2$, this is consistent with the fourth scenario in Fig. 1.